# Step-by-Step Mastery: Enhancing Soft Constraint Following Ability of Large Language Models

### v2

@@ 5 Conclusion
In this paper, we systematically study how to improve LLMs’ ability to follow instructions with soft constraints. Initially, we design a pipeline to automate the construction of datasets with high-quality outputs for soft constraint following. Based on the pipeline, we introduce a method utilizing positive and negative samples generated during the pipeline. Moreover, we propose a new training paradigm that leverages curriculum learning to enhance LLMs’ constraint following ability. Our experiments show that our methods enhance models’ ability to follow soft constraints effectively while maintaining general capabilities.
@@ 6 Limitations
We discuss the limitations of our study as follows. First, we improve the model’s ability to follow complex constraints, thereby improving its overall instruction following capability. However, even when the model’s output meets all the specified constraints, it may still struggle to fully comply with complex instructions due to limitations in reasoning capacity or the knowledge it masters. Additionally, while the dataset constructed in the study encompass a diverse set of tasks, it may still not cover some task types in the long tail. We consider these as a key direction for future research.
Appendix A Details of Data
@@ A.1 Details of Soft Constraints
We utilize LLMs (i.e., GPT-4o) to construct constraints. The three categories of soft constraints that we define are as follows:
@@ A.2 Details of Judger Reordering
Model Reponses Output (a) : Quantum physics studies the behavior of tiny particles. The *wave function* describes a particle’s state, while *superposition state* means a particle can exist in multiple states at once.
@@
You are a helpful assistant who reviews a debate between two other assistants in evaluating the quality of the outputs for a given instruction.The two assistants, Assistant (a) and Assistant (b), are given an instruction, Output (a) and Output (b). They are asked to select the Output (a) or Output (b) that is better for the given instruction. Output (a) and Output (b) are generated by two different AI chatbots respectively. Assistant (a) and Assistant (b) have conflicting evaluations. Your goal is to review their evaluations and give your final decision on which output is better. Here are some rules of the evaluation: (1) You should prioritize evaluating whether the output honestly/precisely/closely executes the instruction, then consider its helpfulness, accuracy, level of detail, harmlessness, etc. (2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction. (3) You should avoid any potential bias and your judgment should be as objective as possible. For example, the order in which the outputs were presented should NOT affect your judgment, as Output (a) and Output (b) are **equally likely** to be the better. Output your final verdict by strictly following this format: "[[A]]" if Output (a) is better, "[[B]]" if Output (b) is better, and "[[C]]" for a tie. /* Given instruction */ {question} /* The Start of Output (a) */ {answer of assistant a} /* The End of Output (a) */ /* The Start of Output (b) */ {answer of assistant b} /* The End of Output (b) */
@@
Appendix B Details of Experiments
B.1 Training hyperparameters
We train Mistral-7B-Instruct-v0.3 and Llama-3-8B-Instruct using LLaMA-Factory Zheng et al. (2024) on 4 NVIDIA A100 80GB GPUs, applying LoRA Hu et al. (2021) for efficient training. The lora target is set to all, and both models use the following training parameters, with training running for 3 epochs. The per device train batch size is set to 1, and gradient accumulation steps is set to 8. The warm-up ratio is set to 0.1. For SFT, Mistral-7B-Instruct-v0.3 is trained with a learning rate 5.0e-7 , while the learning rate of Llama-3-8B-Instruct is 1.0e-4. For DPO, the learning rate is set to 5.0e-6, with a beta value of 0.1.
B.2 Full Results on FollowBench
We present the full results on FollowBench includng the Hard Satisfaction Rate (HSR) metric and Soft Satisfaction Rate (SSR) metric in Tab. B.2.
Model BaseModel FollowBench (HSR) FollowBench (SSR)
L1 L2 L3 L4 L5 Avg L1 L2 L3 L4 L5 Avg
GPT4 Achiam et al. (2023) GPT 84.7 75.6 70.8 73.9 61.9 73.4 84.7 77.0 75.3 77.0 72.3 77.2
GPT3.5-turbo GPT 80.3 68.0 68.6 61.1 53.2 66.2 80.3 71.2 74.2 69.6 67.1 72.5
[…]
Qwen2-72B-Instruct Yang et al. (2024) Qwen 67.9 56.6 47.8 42.2 35.3 50.0 78.0 71.7 69.3 65.1 65.6 69.9
[…]
Conifer-13B Sun et al. (2024) LLaMA2 60.5 53.6 48.4 40.7 31.7 47.0 60.5 58.3 58.2 53.9 51.1 56.4
Vicuna-13B-v1.5 Chiang et al. (2023) LLaMA2 71.2 60.2 49.6 40.6 34.0 51.1 71.2 64.8 59.9 54.5 53.6 60.8
Conifer-7B-SFT Sun et al. (2024) Mistral 54.3 49.5 49.3 40.8 30.5 44.9 53.9 57.6 53.7 54.5 49.7 53.9
Conifer-7B-DPO Sun et al. (2024) Mistral 60.3 53.6 48.0 47.1 41.0 50.0 60.3 55.7 55.7 55.9 53.3 56.2
[…]
Table 14: Full results on FollowBench. We use boldface for the best results and underline for the second-best results among the models ranging from 7B to 13B parameter sizes. indicates that the results are directly sourced from the original benchmarks.

### v3

@@ 5 Conclusion
In this paper, we systematically study how to improve LLMs’ ability to follow soft constraints. Initially, we design a pipeline to automate the construction of datasets with high-quality outputs for soft constraint following. Based on the pipeline, we introduce a method utilizing positive and negative samples generated during the pipeline. Moreover, we propose a new training paradigm that leverages curriculum learning to enhance LLMs’ soft constraint following ability. The experiment results show that our methods enhance models’ ability to follow soft constraints effectively.
@@ 6 Limitations
We discuss the limitations of our study as follows. First, we improve the model’s ability to follow soft constraints, thereby improving its overall instruction following capability. However, even when the model’s output meets all the specified constraints, it may still struggle to fully comply with complex instructions due to limitations in reasoning ability or the knowledge it masters. Additionally, while the dataset constructed encompasses a diverse set of tasks, it may still not cover some task types in the long tail. We consider these as key directions for future research.
Appendix A Appendix
@@ A.1 Details of Soft Constraints
We utilize GPT-4o to construct soft constraints. The three categories of soft constraints that we define are as follows:
@@ A.2 Details of Judger Reordering
Model Responses Output (a) : Quantum physics studies the behavior of tiny particles. The *wave function* describes a particle’s state, while *superposition state* means a particle can exist in multiple states at once.
@@
You are a helpful assistant who reviews a debate between two other assistants in evaluating the quality of the outputs for a given instruction.The two assistants, Assistant (a) and Assistant (b), are given an instruction. Output (a) and Output (b) are generated by two different AI chatbots respectively. Assistant (a) and Assistant (b) have conflicting evaluations. Your goal is to review their evaluations and give your final decision on which output is better. Here are some rules of the evaluation: (1) You should prioritize evaluating whether the output honestly/precisely/closely executes the instruction, then consider its helpfulness, accuracy, level of detail, harmlessness, etc. (2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction. (3) You should avoid any potential bias and your judgment should be as objective as possible. For example, the order in which the outputs were presented should NOT affect your judgment, as Output (a) and Output (b) are **equally likely** to be the better. Output your final verdict by strictly following this format: "[[A]]" if Output (a) is better, "[[B]]" if Output (b) is better, and "[[C]]" for a tie. /* Given instruction */ {question} /* The Start of Output (a) */ {answer of assistant a} /* The End of Output (a) */ /* The Start of Output (b) */ {answer of assistant b} /* The End of Output (b) */
@@
A.3 Implementation Details
We train LLaMA-3-8B-Instruct, Mistral-7B-Instruct-v0.3, and LLaMA2-13B-Chat-HF using LLaMA-Factory Zheng et al. (2024) on 4 NVIDIA A100 80GB GPUs, applying LoRA Hu et al. (2021) for efficient training. The lora target is set to all, with all models training for 3 epochs. The per device train batch size is set to 1, and gradient accumulation steps is set to 8. The warm-up ratio is set to 0.1. For SFT, LLaMA-3-8B-Instruct has a learning rate of 1.0e-4, Mistral-7B-Instruct-v0.3 uses 5.0e-7, and LLaMA2-13B-Chat-HF uses 1.0e-7. For DPO, the learning rate is 5.0e-6 with a beta value of 0.1. We apply cosine learning rate scheduler.